Friday 02/09/12

10:10 A — 11:20 P

  • move samples from 4C, change 100% EtOH
  • STORM: setup 10x objective for finding samples. substantial offset, need to calibrate.
  • Need to set laser powers to zero but have lasers on before starting dave!
  • Ordered Bloomington GFP stocks:
  • 9593 w[1118]; P{w[+mC]=Pc[T:Avic\GFP-EGFP]}3
    29018 P{w[+mC]=EGFP-bcd}1, y[1] w[1118]; th[1] st[1] kni[ri-1] bcd[6] rn[roe-1] p[p]
    30870 y[1] w[*]; PBac{y[+mDint2] w[+mC]=en-EGFP.S}VK00033
    30871 y[1] w[*]; PBac{y[+mDint2] w[+mC]=eve-EGFP.S}VK00033
    30872 y[1] w[*]; PBac{y[+mDint2] w[+mC]=slp2-EGFP.S}VK00033/TM3, Sb[1]
    30874 y[1] w[*]; PBac{y[+mDint2] w[+mC]=tll-EGFP.S}VK00037
    31419 w[*]; P{w[+mC]=UASp-bnb[N20].mEos}1
  • continuing embryo staining, day 4: embedding.
  • Flip fly cage 1P.  (weak egg lay).
  • STORM: slide H3, 750 might be okay.
  • STORM: slide S2, running O/N.
  • Confocal imaging (see summary posts).
  • Embedding embryos en mass in cut, circular embedding cases.
This entry was posted in Summaries and tagged , , . Bookmark the permalink.